To obtain Grade A, Ben must achieve an average of at least 70 in five tests. if his average mark the first four test is 68, what is the lowest he can get in his fifth test and still obtain Grade A?
1
Expert's answer
2019-07-10T11:46:14-0400

Let a1 be the mark for the first test, a2 be the mark for the second test and so on up to a5. For a5 to be the smallest, the average of all scores must be 70: (a1+a2+a3+a4+a5)/5=70.Because the average of the first four estimates 68: (a1+a2+a3+a4)/4=68 => a1+a2+a3+a4=272 substitute the sum in the previous equation: (272

+a5)/5=70 => 272+a5=350 => a5=78. Answer: 78